What is the dimensions of a rectangle if the length of the rectangle is 2 more than twice its width and the perimeter of the rectangle is 34 in.?
Answer by faceoff57(108)

2L + 2W = 34
2(2+2W) + 2W = 34
4 + 4W + 2W = 34
4 + 6W = 34
6W = 30
W = 5
THE WIDTH IS 5 INCHES.
THE LENGTH IS 12 INCHES.